OpenStudy (anonymous):

given the center (h,k) and radius r=3, write an equation for the circle in standard form a) (x+5)^2+y^2=9 b) (x-5)^2+y^2=9 c) (x-5)^2+y^2=3

OpenStudy (solomonzelman):

The equation of a circle is: \(\large\color{black}{ \displaystyle (x-{\rm \color{red}{h}} )^2+ (y-{\rm \color{blue}{k}} )^2 = {\rm \color{green}{r}}^2 }\) Where 1) \(\large\color{black}{ \displaystyle({\rm \color{red}{h}} ,~{\rm \color{blue}{k}} )}\) is the \(\normalsize\color{red}{ \bf c\color{blue}{e}n\color{blue}{t}e\color{blue}{r} }\) of the circle. 2) \(\large\color{black}{ \displaystyle {\rm \color{green}{r}} }\) is the \(\normalsize\color{green}{ \bf radius }\) of the circle.

OpenStudy (solomonzelman):

this is the general form, and you don't even need to change the (h,k). All you need to do is to substitute 3 for \(\rm \color{green}{r}\).

OpenStudy (solomonzelman):

you have some typos in a question apparently....

OpenStudy (anonymous):

its B ? i forgot to put that (h,k)=(5,0)

OpenStudy (solomonzelman):

ok, h=5 k=0 r=3 then, yes, option B is correct. (x-5)²+y²=9

OpenStudy (solomonzelman):

questions?

OpenStudy (anonymous):

no thank you !

OpenStudy (solomonzelman):

Alright! You are welcome; have a good luck!